OpenStudy (anonymous):

Solve by factoring the trinomial as a binomial square and then use the square root method. Round to the nearest tenth if necessary. x2 + 16x + 64 = 81 A. {9, -9} B. {-8, 9} C. {-17, 1} D. no real solution Solve by factoring the trinomial as a binomial square and then use the square root method. Round to the nearest tenth if necessary. x2 + 16x + 64 = 81 A. {9, -9} B. {-8, 9} C. {-17, 1} D. no real solution @Mathematics

hero (hero):

(x+8)^2 = 81 x+8 = 9 x = 1

hero (hero):

and -17

hero (hero):

C

OpenStudy (anonymous):

x2 + 16x + 64 = 81 (x)^2 + 2(x)(8) + (8)^2 = 81 (x + 8)^2 = 81 taking root of both sides x + 8 = ± sqroot(81) x + 8 = ± 9 x = -8 ± 9 x = -8 - 9 or x = -8 + 9 x = -17 or x = 1

OpenStudy (anonymous):

so correct solution is C. {-17, 1}